OpenStudy (anonymous):

write the slope-intercept equation for the line containing the given pair of points (-3,4) and (3,1)

Parth (parthkohli):

Kill me if this doesn't help :P http://openstudy.com/study#/updates/4fb9f5a9e4b05565342ee28c

OpenStudy (anonymous):

what then is the equation that pass through the given ordered pairs? @ParthKohli

Parth (parthkohli):

First find the slope, and then the y-intercept.

Parth (parthkohli):

You can find the slope with its formula. \(\Large \color{MidnightBlue}{\Rightarrow {1 - 4 \over 3 + 3} }\) \(\Large \color{MidnightBlue}{\Rightarrow {-3 \over 6} = -{1 \over 2} }\)

Parth (parthkohli):

Then, the standard equation is y = mx + b Slope is m, and it is -1/2. \(\Large \color{MidnightBlue}{\Rightarrow y = {-1 \over 2}x + b }\) Take any pair of points (x,y) given and substitute to find b.

Parth (parthkohli):

Have you gotten the hint till now?

OpenStudy (anonymous):

m=1-(4)/3-(-3) m=-3/3-(-3) m=-3/3+3 m=-3/6 m=-1/2

OpenStudy (anonymous):

yes I have got it

OpenStudy (anonymous):

Thank you